OpenStudy (anonymous):

Find the zeros of the polynomial function with equation f(x) = x3-x2 -30x x = 0, x = 4, x = 5 x = 4, x = 5 x = -6, x = 5 x = 0, x = 6, x = -5

OpenStudy (anonymous):

0, 6, and -5

OpenStudy (cadteach):

Factoring out an x, f(x) = x (x^2-x-30). so x = 0 is one. THen factor x^2 - x -30, to be (x-6)(x+5), so the others are 6 and -5.

OpenStudy (babyslapmafro):

f(x)=x3-x2-30=x(x2-x-30)=x(x-6)(x+5)=0